The perimeter of a rectangle is 482cm. The length is 6cm greater than 4 times the width. Find the length

Answer by larryblackwell(1)

L=6+4x,w=x,~P=2L+2W~ 482=2[6+4x]+2x~ 482=12+8x+2x ~ 482=12+10x ~ 482-12=12-12+10x ~470=10x ~ 470/10=10x/10 ~ 47=x : L=6+4[47]~ L=6+188 therefore L=194cm.